# Counterfactual Reasoning for Multi-Label Image Classification via Patching-Based Training

### v1

@@ Abstract
The key to multi-label image classification (MLC) is to improve model performance by leveraging label correlations. Unfortunately, it has been shown that overemphasizing co-occurrence relationships can cause the overfitting issue of the model, ultimately leading to performance degradation. In this paper, we provide a causal inference framework to show that the correlative features caused by the target object and its co-occurring objects can be regarded as a mediator, which has both positive and negative impacts on model predictions. On the positive side, the mediator enhances the recognition performance of the model by capturing co-occurrence relationships; on the negative side, it has the harmful causal effect that causes the model to make an incorrect prediction for the target object, even when only co-occurring objects are present in an image. To address this problem, we propose a counterfactual reasoning method to measure the total direct effect, achieved by enhancing the direct effect caused only by the target object. Due to the unknown location of the target object, we propose patching-based training and inference to accomplish this goal, which divides an image into multiple patches and identifies the pivot patch that contains the target object. Experimental results on multiple benchmark datasets with diverse configurations validate that the proposed method can achieve state-of-the-art performance.
[…]
@@ 2 Related Work
With the rapid development of deep models, MLC has recently attracted increasing attention. Existing solutions to the MLC problems can be roughly divided into three groups. The first kind of methods improved the traditional BCE loss to address the inherent positive-negative imbalance problem in the MLC tasks (Ridnik et al., 2021b). The second kind focused on identifying regions of interest associated with semantic labels by utilizing the global average pooling (GAP) (Verelst et al., 2023) strategy or the attention technique (Lanchantin et al., 2021; Liu et al., 2021a; Ridnik et al., 2023b). The last kind aimed at modeling label correlations, which have been regarded as fundamental information for MLC. The representative methods adopted a GCN (Chen et al., 2019b) or developed a label masking strategy (Lanchantin et al., 2021) to explore label correlations during training.
With the intensive consideration of correlations among variables, Peal’s causal graphs (Pearl, 2009, 2001) and causal inference methods are widely used in machine learning. It provides solutions for achieving goals, e.g., improving model robustness by pursuing causal effects. Intervention-based and counterfactual-inference-based approaches were primarily utilized to achieve these goals. The intervention is mainly reflected in the control of confounding variables. For example, VC R-CNN (Wang et al., 2020a) applied causal intervention to unsupervised regional feature learning, aiming to introduce a kind of “visual common sense” into model design. CONTA (Zhang et al., 2020) attributed the cause of ambiguous pseudo-mask boundaries to the confusion context, and then removed the confusion bias via the backdoor criterion for weakly supervised semantic segmentation. CaaM (Wang et al., 2021), as a novel causal attention module, self-annotated confounders in an unsupervised method and decoupled confounders and mediators via adversarial training. On the other hand, counterfactual-based techniques have often been applied to control or analyze causal relationships for tasks of interest. For example, while momentum may inadvertently bias tail predictions toward head predictions when training on long-tailed data, it also enhances representation learning and head predictions through an induced mediator (Tang et al., 2020). This paradoxical effect of momentum could be effectively disentangled by optimizing the total direct effect (TDE).
@@ 3.1 Notations
(a) Conditional True Positive Ratio
(b) Conditional False Positive Ratio
@@ 4 The Proposed Method
To obtain the predicted probability $P(Y^{k}_{\bm{o}}=1|X=\bm{x},Z=\bm{z}_{0})$, we propose a patching-based inference (PAT-I) method. The main idea involves partitioning the original image into multiple patches and subsequently determining which patch contains the target object. Then, the prediction for the identified patch can be considered as an approximation of the prediction for the target object. Specifically, during inference, for each image $\bm{i}$, we crop it into $m$ equally-sized patches $\{\bm{u}_{j}\}_{j=1}^{m}$. By feeding the original image and the patches into the trained model $f(\cdot)$, we obtain the predicted probability $\bm{p}=[p_{1},...,p_{q}]$ and $\bm{q}_{j}=[q_{j1},...,q_{jq}]$, where $p_{k}=f_{k}(\bm{i})$ and $q_{jk}=f_{k}(\bm{u}_{j})$ denote the predicted probability of the class $k$ for the image $\bm{i}$ and patch $\bm{u}_{j}$ respectively.
Given the probability that the target object may be present in multiple patches, rather than selecting a particular patch, we assign a weight to each patch based on its prediction
@@
where $\tau$ is a temperature parameter that controls the degree of emphasis on a particular patch. Then, the predicted probabilities for different objects can be obtained by aggregating the probabilities $\bm{q}_{j}$ in a spatially-weighted manner, $\bm{q}=\sum\limits_{j=1}^{m}\bm{w}_{j}\cdot\bm{q}_{j}$. Finally, according to Eq. (3), we obtain the model prediction for $\bm{x}$ of class $k$ as $\text{TDE}_{k}(\bm{i})=p_{k}+\lambda\cdot q_{k}$. In our experiments, we observed that setting $\lambda=1$ achieves favorable performance. The detailed procedures of PAT-I are summarized in Appendix C. Generally, PAT-I has two main advantages: i) simplicity - it can be incorporated directly into the inference process to enhance model performance without any additional training; ii) flexibility - it can be seamlessly integrated into any trained model.
@@
Specifically, we decompose the trained model $f(\cdot)$ into two components $f(\cdot)=h\circ g(\cdot)$, where $g(\cdot)$ is a backbone and $h(\cdot)$ is a classification head. During training, for each image $\bm{i}$, we obtain the predictions $\bm{p}$ and $\bm{q}_{j}$ in the same way. In practice, to further improve the model performance, we make two modifications. Firstly, considering that the classification head is sensitive to two different classification tasks on images and patches, we adopt two classifier heads $h_{\phi}$ and $h_{\psi}$ to make predictions on images and patches independently, i.e., $\bm{p}=h_{\phi}\circ g(\bm{i})$ and $\bm{q}_{j}=h_{\psi}\circ g(\bm{u}_{j})$, where $\phi$ and $\psi$ denote two sets of parameters. Secondly, we also use another head $h_{\theta}$ to obtain the predicted probabilities $\bm{q}^{w}_{j}=h_{\theta}\circ g(\bm{u}_{j})$, which are used to estimate weights $\bm{w}_{j}$ according to Eq. (4). By decoupling the dependence between weight estimation and probability generation, it enhances model performance by avoiding the accumulation of prediction errors on a single head.
@@
$\mathcal{L}=\ell(\bm{p},\bm{y})+\ell(\bm{q},\bm{y}),$ (5)
where $\ell(\cdot,\cdot)$ can be BCE or ASL loss. In our experiments, we use ASL loss because it often achieves better performance than BCE loss. The detailed procedures of PAT-T are presented in Appendix C.
@@ 5.1 Experimental Settings
Table 1: Comparison results on COCO between our method and state-of-the-art methods. The backbones denoted with 22k and OI are pretrained on the ImageNet-22k and Open Image datasets, respectively. The best performance is highlighted in bold. All metrics are in %.
@@
KSSNet (Wang et al., 2020b) ResNet101 $448\times 448$ 83.7 84.6 73.2 77.2 87.8 76.2 81.5
@@
To evaluate the performance of the proposed method, we conduct experiments on three benchmark datasets, including MS-COCO 2014  [^2] https://cocodataset.org (Lin et al., 2014), Pascal VOC 2007  [^3] http://host.robots.ox.ac.uk/pascal/VOC/ (Everingham et al., 2010), and Visual Genome  [^4] https://homes.cs.washington.edu/~ranjay/visualgenome/index.html (Krishna et al., 2017). MS-COCO contains 82,081 training images and 40,137 validation images for 80 classes, with an average of 2.9 labels per image. VOC 2007 consists of the train-val set with 5,011 images and test set with 4,952 images. Following previous works (Ridnik et al., 2021b; Liu et al., 2021a), we use train-val set for training and test set for testing. Visual Genome is a dataset that contains 108,249 images and covers 80,138 categories. Similar to the previous work (Chen et al., 2022), considering that most categories have very few examples, we conduct preprocessing on this dataset. The main steps include: i) merge categories: combine categories that share the same meaning; ii) filter by image count: exclude categories with fewer than 500 images. Finally, we obtain a dataset named VG-200 that contains 201 classes and 108,249 images.
@@
Following previous works (Ridnik et al., 2021b; Liu et al., 2021a), we choose TResNetL (Ridnik et al., 2021c) pre-trained on ImageNet (Deng et al., 2009) or Open Image (Papadopoulos et al., 2017) as the backbone for our model, as it demonstrates superior performance compared to the conventional ResNet-101 (He et al., 2016) in the context of MLC while adhering to similar efficiency constraints on GPU. We employ Adam (Kingma & Ba, 2014) optimizer and one-cycle policy scheduler to train the model with maximal learning rate of 0.0001. Furthermore, we perform exponential moving average (EMA) (Tarvainen & Valpola, 2017) for the model parameters with a decay of 0.9997. We use the ASL loss as the base loss function since it shows superiority to BCE loss. We perform all experiments on GeForce RTX 3090 GPUs. The random seed is set to 1 for all experiments.
@@ 6 Comparison with State-of-the-art Methods
6.1 Performance on MS-COCO
Table 2: Comparison results on VOC between our method and state-of-the-art methods in terms of AP (%) and mAP (%). All results are reported at resolution $448\times 448$ except for the ADD-GCN and SSGRL, whose resolutions are $576\times 576$. The best performance is highlighted in bold.
@@ 6.1 Performance on MS-COCO
Table 3: Comparison results on VG-200 between our method and state-of-the-art methods. All results are reported at resolution $448\times 448$. The best performance for each backbone is highlighted in bold.
@@
ResNet101 ResNet101 52.5 54.9 52.7 53.8 62.1 63.9 63.0
PAT-I ResNet101 53.1 57.0 51.4 54.0 62.2 65.2 63.7
PAT-T ResNet101 54.8 55.5 56.2 55.8 62.4 65.8 64.1
TResNetL TResNetL (22k) 58.3 60.3 57.7 59.0 65.2 67.1 66.1
PAT-I TResNetL (22k) 59.2 61.6 57.4 59.4 64.0 69.1 66.4
PAT-T TResNetL (22k) 60.8 61.6 60.6 61.1 65.9 68.8 67.3
Q2L-TResL TResNetL (22k) 60.6 59.4 63.1 61.2 64.6 69.5 67.0
PAT-I+Q2L-TResL TResNetL (22k) 61.2 58.9 63.8 61.2 62.1 72.1 66.7
PAT-T+Q2L-TResL TResNetL (22k) 62.9 60.7 64.2 62.4 66.1 69.9 67.9
@@ 6.2 Performance on VOC 2007
Table 2 reports the comparison results on VOC 2007. To make a fair comparison, following the previous works (Ridnik et al., 2021b; Liu et al., 2021a), we use TResNet pre-trained on ImageNet-1K as the backbone. The results based on a stronger backbone can be found in Appendix B. Our method achieves the best performance when compared to state-of-the-art methods. Moreover, our method achieves the optimal performance in the majority of classes. It seems that our method does not yield as significant improvement as it did on COCO. Except for the small number of images and relatively high performance achieved on VOC, one possible reason may be that the co-occurrence relationships in the dataset are not as strong as in COCO, as each example has only 1.5 label on average. This makes models less prone to overfit to label correlations.
6.3 Performance on VG-200
Table 3 reports the comparison results between our methods PAT-I, PAT-T and the comparing methods on VG-200. We run all methods in the resolution of $448\times 448$. From the table, we can see that PAT-I can improve the model performance without any additional training cost. It enhances the positive impact while alleviating the negative impact of the mediated effect during the inference phase. Furthermore, by incorporating this idea into training, PAT-T achieves significant performance improvements and obtains the best performance. These results strongly verify the effectiveness of the proposed method.
@@ 6.4 Ablation Studies
To demonstrate the effectiveness of our method in preventing models from overfitting to co-occurrence relationships. Figure 5 respectively illustrates the conditional FPR and conditional TPR of label pairs with a co-occurrence probability greater than 0.2. Baseline (ASL) and our PAT-T are trained with the same configuration (the backbone is TResNetL). In order to reduce the apparent fluctuations in the curve, we sort the label pairs according to the performance of Baseline. From the figure, we can see that the curve of PAT-T is generally below ASL, yielding that PAT-T can prevent the model from overfitting to co-occurrence, thereby making fewer mistakes in predicting false positive labels. Moreover, PAT-T also outperforms Baseline in terms of conditional TPR, indicating that PAT-T enhances the recognition capacity of the model significantly.

### v2

@@ Abstract
The key to multi-label image classification (MLC) is to improve model performance by leveraging label correlations. Unfortunately, it has been shown that overemphasizing co-occurrence relationships can cause the overfitting issue of the model, ultimately leading to performance degradation. In this paper, we provide a causal inference framework to show that the correlative features caused by the target object and its co-occurring objects can be regarded as a mediator, which has both positive and negative impacts on model predictions. On the positive side, the mediator enhances the recognition performance of the model by capturing co-occurrence relationships; on the negative side, it has the harmful causal effect that causes the model to make an incorrect prediction for the target object, even when only co-occurring objects are present in an image. To address this problem, we propose a counterfactual reasoning method to measure the total direct effect, achieved by enhancing the direct effect caused only by the target object. Due to the unknown location of the target object, we propose patching-based training and inference to accomplish this goal, which divides an image into multiple patches and identifies the pivot patch that contains the target object. Experimental results on multiple benchmark datasets with diverse configurations validate that the proposed method can achieve state-of-the-art performance. The implementation is available at https://github.com/xiemk/MLC-PAT.
Machine Learning, ICML
@@ 2 Related Work
With the rapid development of deep models, MLC has recently attracted increasing attention. Existing solutions to the MLC problems can be roughly divided into three groups. The first kind of methods improved the traditional BCE loss to address the inherent positive-negative imbalance problem in the MLC tasks (Ridnik et al., 2021b). The second kind focused on identifying regions of interest associated with semantic labels by utilizing the global average pooling (GAP) (Verelst et al., 2023) strategy or the attention technique (Lanchantin et al., 2021; Liu et al., 2021a; Ridnik et al., 2023b). The last kind aimed at modeling label correlations, which have been regarded as fundamental information for MLC. The representative methods adopted a GCN (Chen et al., 2019b) or developed a label masking strategy (Lanchantin et al., 2021) to explore label correlations during training. Recent works study how to improve multi-label classification performance with limited supervision, including semi-supervised multi-label learning (Xie et al., 2023; Wang et al., 2020a), multi-label learning with partial labels (Durand et al., 2019; Xie et al., 2022), and multi-label learning with noisy labels (Xie & Huang, 2022; Li et al., 2022).
With the intensive consideration of correlations among variables, Peal’s causal graphs (Pearl, 2009, 2001) and causal inference methods are widely used in machine learning. It provides solutions for achieving goals, e.g., improving model robustness by pursuing causal effects. Intervention-based and counterfactual-inference-based approaches were primarily utilized to achieve these goals. The intervention is mainly reflected in the control of confounding variables. For example, VC R-CNN (Wang et al., 2020b) applied causal intervention to unsupervised regional feature learning, aiming to introduce a kind of “visual common sense” into model design. CONTA (Zhang et al., 2020) attributed the cause of ambiguous pseudo-mask boundaries to the confusion context, and then removed the confusion bias via the backdoor criterion for weakly supervised semantic segmentation. CaaM (Wang et al., 2021), as a novel causal attention module, self-annotated confounders in an unsupervised method and decoupled confounders and mediators via adversarial training. On the other hand, counterfactual-based techniques have often been applied to control or analyze causal relationships for tasks of interest. For example, while momentum may inadvertently bias tail predictions toward head predictions when training on long-tailed data, it also enhances representation learning and head predictions through an induced mediator (Tang et al., 2020). This paradoxical effect of momentum could be effectively disentangled by optimizing the total direct effect (TDE).
(a) Conditional True Positive Ratio
(b) Conditional False Positive Ratio
@@ 4 The Proposed Method
To obtain the predicted probability $P(Y^{k}_{\bm{o}}=1|X=\bm{x},Z=\bm{z}_{0})$, we propose a patching-based inference (PAT-I) method. The main idea involves partitioning the original image into multiple patches and subsequently determining which patch contains the target object. Then, the prediction for the identified patch can be considered as an approximation of the prediction for the target object. Specifically, during inference, for each image $\bm{i}$, we crop it into $m$ equally-sized patches $\{\bm{u}_{j}\}_{j=1}^{m}$. By feeding the original image and the patches into the trained model $f(\cdot)$, we obtain the predictions $\bm{p}=[p_{1},...,p_{q}]$ and $\bm{q}_{j}=[q_{j1},...,q_{jq}]$, where $p_{k}=f_{k}(\bm{i})$ and $q_{jk}=f_{k}(\bm{u}_{j})$ denote the predicted logits of the class $k$ for the image $\bm{i}$ and patch $\bm{u}_{j}$ respectively.
Given the probability that the target object may be present in multiple patches, rather than selecting a particular patch, we assign a weight to each patch based on its predictions
@@
where $\tau$ is a temperature parameter that controls the degree of emphasis on a particular patch. Then, the predicted logits for different objects can be obtained by aggregating the logits $\bm{q}_{j}$ in a spatially-weighted manner, $\bm{q}=\sum\limits_{j=1}^{m}\bm{w}_{j}\cdot\bm{q}_{j}$. Finally, according to Eq. (3), we obtain the model prediction for $\bm{x}$ of class $k$ as $\text{TDE}_{k}(\bm{i})=\delta(p_{k}+\lambda\cdot q_{k}$), where $\delta(\cdot)$ is the Sigmoid function. In our experiments, we observed that setting $\lambda=1$ achieves favorable performance. The reason for performing above operations on logits instead of probabilities is that we found applying the Sigmoid function to logits to convert them into probabilities compresses the distinctive information, which in turn hinders the effectiveness of our method. Generally, PAT-I has two main advantages: i) simplicity - it can be incorporated directly into the inference process to enhance model performance without any additional training; ii) flexibility - it can be seamlessly integrated into any trained model.
@@
Specifically, we decompose the trained model $f(\cdot)$ into two components $f(\cdot)=h\circ g(\cdot)$, where $g(\cdot)$ is a backbone and $h(\cdot)$ is a classification head. During training, for each image $\bm{i}$, we obtain the predictions $\bm{p}$ and $\bm{q}_{j}$ in the same way. In practice, to further improve the model performance, we make two modifications. Firstly, considering that the classification head is sensitive to two different classification tasks on images and patches, we adopt two classifier heads $h_{\phi}$ and $h_{\psi}$ to make predictions on images and patches independently, i.e., $\bm{p}=h_{\phi}\circ g(\bm{i})$ and $\bm{q}_{j}=h_{\psi}\circ g(\bm{u}_{j})$, where $\phi$ and $\psi$ denote two sets of parameters. Secondly, we also use another head $h_{\theta}$ to obtain the predicted logits $\bm{q}^{w}_{j}=h_{\theta}\circ g(\bm{u}_{j})$, which are used to estimate weights $\bm{w}_{j}$ according to Eq. (4). By decoupling the dependence between weight estimation and logits prediction, it enhances model performance by avoiding the accumulation of prediction errors on a single head.
@@
$\mathcal{L}=\ell(\delta(\bm{p}),\bm{y})+\ell(\delta(\bm{q}),\bm{y}),$ (5)
where $\ell(\cdot,\cdot)$ can be BCE or ASL loss. In our experiments, we use ASL loss because it often achieves better performance than BCE loss.
@@ 5.1 Experimental Settings
Table 1: Comparison results on MS-COCO between our method and state-of-the-art methods. The backbones denoted with 21k and OI are pretrained on the ImageNet-21k and Open Image datasets, respectively. The best performance is highlighted in bold. All metrics are in %.
@@
KSSNet (Wang et al., 2020c) ResNet101 $448\times 448$ 83.7 84.6 73.2 77.2 87.8 76.2 81.5
@@
To evaluate the performance of the proposed method, we conduct experiments on three benchmark datasets, including MS-COCO 2014  [^2] https://cocodataset.org (Lin et al., 2014), Pascal VOC 2007  [^3] http://host.robots.ox.ac.uk/pascal/VOC/ (Everingham et al., 2010), and Visual Genome  [^4] https://homes.cs.washington.edu/~ranjay/visualgenome/index.html (Krishna et al., 2017). MS-COCO contains 82,081 training images and 40,137 validation images for 80 classes, with an average of 2.9 labels per image. VOC 2007 consists of the train-val set with 5,011 images and test set with 4,952 images. Following previous works (Ridnik et al., 2021b; Liu et al., 2021a), we use train-val set for training and test set for testing. Visual Genome is a dataset that contains 108,249 images and covers 80,138 categories. Similar to the previous work (Chen et al., 2022), considering that most categories have very few examples, we conduct preprocessing on this dataset. The main steps include: i) merge categories: combine categories that share the same meaning; ii) filter by image count: exclude categories with fewer than 500 images. Finally, we obtain a dataset named VG-256 that contains 256 classes and 108,249 images. We randomly divide 70% for training and 30% for testing.
@@
Following previous works (Ridnik et al., 2021b; Liu et al., 2021a), we choose ResNet101 and TResNetL (Ridnik et al., 2021c) pre-trained on ImageNet (Deng et al., 2009) or Open Image (Papadopoulos et al., 2017) as the backbone for our method. We employ Adam (Kingma & Ba, 2014) optimizer and one-cycle policy scheduler to train the model with maximal learning rate of 0.0001. Furthermore, we perform exponential moving average (EMA) (Tarvainen & Valpola, 2017) for the model parameters with a decay of 0.9997. We use the ASL loss as the base loss function since it shows superiority to BCE loss. We perform all experiments on GeForce RTX 3090 GPUs. The random seed is set to 1 for all experiments.
@@ 6 Comparison with State-of-the-art Methods
Table 2: Comparison results on VOC 2007 between our method and state-of-the-art methods in terms of AP (%) and mAP (%). All results are reported at resolution $448\times 448$ except for the ADD-GCN and SSGRL, whose resolutions are $576\times 576$. The best performance is highlighted in bold.
@@
Table 3: Comparison results on VG-256 between our method and state-of-the-art methods. All results are reported at resolution $448\times 448$. The best performance for each backbone is highlighted in bold.
@@
ResNet101 ResNet101 47.7 51.1 48.2 49.6 63.3 64.0 63.6
PAT-I ResNet101 48.5 53.9 46.3 49.8 64.0 64.8 64.4
PAT-T ResNet101 49.9 52.0 51.7 51.9 63.5 66.0 64.7
TResNetL TResNetL (21k) 53.9 57.2 53.4 55.2 66.7 67.7 67.2
PAT-I TResNetL (21k) 54.8 58.7 52.5 55.4 66.1 69.1 67.6
PAT-T TResNetL (21k) 56.2 58.0 56.0 57.0 67.1 69.3 68.2
Q2L-TResL TResNetL (OI) 57.7 58.7 58.6 58.7 67.4 70.6 69.0
PAT-I+Q2L-TResL TResNetL (OI) 58.0 58.8 58.3 58.6 66.1 72.0 68.9
PAT-T+Q2L-TResL TResNetL (OI) 59.5 59.6 59.9 59.7 67.9 71.7 69.8
6.1 Performance on MS-COCO
@@ 6.2 Performance on VOC 2007
Table 2 reports the comparison results on VOC 2007. To make a fair comparison, following the previous works (Ridnik et al., 2021b; Liu et al., 2021a), we use TResNet pre-trained on ImageNet-1K as the backbone. The results based on a stronger backbone can be found in Appendix B. Our method achieves the best performance when compared to state-of-the-art methods. Moreover, our method achieves the optimal performance in the majority of classes. It seems that our method does not yield as significant improvement as it did on MS-COCO. Except for the small number of images and relatively high performance achieved on VOC 2007, one possible reason may be that the co-occurrence relationships in the dataset are not as strong as in MS-COCO, as each example has only 1.5 labels on average. This makes models less prone to overfitting to label correlations.
6.3 Performance on VG-256
Table 3 reports the comparison results between our methods PAT-I, PAT-T and the comparing methods on VG-256. We run all methods in the resolution of $448\times 448$. From the table, we can see that PAT-I can improve the model performance without any additional training cost. It enhances the positive impact while alleviating the negative impact of the mediated effect during the inference phase. Furthermore, by incorporating this idea into training, PAT-T achieves significant performance improvements and obtains the best performance. These results strongly verify the effectiveness of the proposed method.
Figure 5: Conditional FPR and conditional TPR of the label pairs with co-occurrence probabilities larger than 0.2 on MS-COCO. The pair indices are sorted according to the performance of Baseline. PAT-T outperforms Baseline in around 76% and 77% percentage of label pairs in terms of CFPR and CTPR.
@@ 6.4 Ablation Studies
To demonstrate the effectiveness of our method in preventing models from overfitting to co-occurrence relationships, Figure 5 respectively illustrates the conditional FPR and conditional TPR of label pairs with a co-occurrence probability greater than 0.2. Baseline (ASL) and our PAT-T are trained with the same configuration (the backbone is TResNetL). In order to reduce the apparent fluctuations in the curve, we sort the label pairs according to the performance of Baseline. From the figure, we can see that the curve of PAT-T is generally below ASL, yielding that PAT-T can prevent the model from overfitting to co-occurrence, thereby making fewer mistakes in predicting false positive labels. Moreover, PAT-T also outperforms Baseline in terms of conditional TPR, indicating that PAT-T enhances the recognition capacity of the model significantly.